Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Simple Probability Problem. Help with concept needed :-)

Expert replies
by amazonviper » Sat May 16, 2009 6:05 pm
Problem:
A dice is rolled 6 times. What is the Probability that "6" will NOT be facing upward on all three rolls ?.

My method:

Prob that we will not get a "6" first time = 5/6
Prob that we will not get a "6" 2nd time = 5/6
Prob that we will not get a "6" 3rd time = 5/6

So, Prob that we will not get "6" all three times is 5/6*5/6*5/6

Seems that this is incorrect.The book calculates Probability of 6 appearing all three times and then subtracting from 1 to get the answer.Why cannot we use the method above(My method)?

To me, this type of question seems straightforward and I want to be sure that I am not using the wrong logic here

Thanks in advance for the help :-)

Viper
Join the discussion
Source: — Problem Solving |

Hello!

by amazonviper » Mon May 18, 2009 1:44 pm
Any help is greatly appreciated!
Join the discussion

by ManSab » Mon May 18, 2009 3:34 pm
total outcomes in 3 rolls = 18
6 can face up 3 times in 3 rolls.

probability for 6 facing up in 3 rolls = 3/18 = 1/6
probability 6 will not face = 1- 1/6 = 5/6.

Please correct if I am wrong.

Thanks
Join the discussion

by dumb.doofus » Tue May 19, 2009 3:57 pm
ok.. well here's the problem with your solution.. and first of all thanks for asking this question.. quite a conceptual topic..

when you say that the probability of not getting a 6 is 5/6, the first time and second time and third time.. essentially the you are finding all combinations where 6 is not used at all.. examples are:

1 2 3
3 4 5
4 5 2
1 3 5
5 3 1 and I could go on and on 125 times.. that is what your answer is...

But this answer doesn't take into account the following choices:

1 3 6
6 4 5
6 4 2
3 6 5 and so on.. 90 of them to be precise.

Now the reason why your solution is not taking into account these choices.. the problem is that you have already mentioned that there are 5 ways to choose for first number, 2nd number and 3rd number.. and that's the reason you are getting 125 ways.. we also have to exclude the above 90 ways tooo

So the correct solution is:

we know that there is only one combination out of 216 combinations, where we will get 6 on all the three rolls.. i.e. 6 6 6

so the probability of getting a 6 on all three is 1/216

so the probability of not getting a 6 at all is 1 - 1/216 = 215/216


Hope this clarifies..
One love, one blood, one life. You got to do what you should.
https://dreambigdreamhigh.blocked/
https://gmattoughies.blocked/
Join the discussion

by mike22629 » Wed May 20, 2009 11:26 am
The 1-x probability is the best way to answer this problem.

Learn how to do this because it will make probability problems on the GMAT a lot easier.

First, find total possibilities of rolls.

6^3 = 216

First roll = 6 possibilities
Second roll = 6 possibilites
Third Roll = 6 possibilites

Now find the total number of possibilities of rolling 6 all three times.

Obviously the answer is 1

So....
216-1 = 215 (possibilities of not rolling 6 all three times)

Answer 215/216

Trust me when I tell you that this is VERY COMMON in probability problems on GMAT. It is worth your time to master the "1-x" probability approach.
Join the discussion